The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 58 September 17, 2001

We extend our deepest sympathies to the victims of the terrorist attacks on September 11.

Contents

Abstracts of 7 refereed papers  1
Thesis abstract  5
Abstracts of 4 non-refereed papers 6

Abstracts of Refereed Papers

The effects of dust in simple environments: Large Magellanic Cloud H ii regions

E. F. Bell^{1}, K. D. Gordon^{1}, R. C. Kennicutt Jr.^{1} and D. Zaritsky^{1}

1 Steward Observatory, 933 N. Cherry Ave., Tucson, AZ 85721, USA

We investigate the effects of dust on Large Magellanic Cloud (LMC) H ii region spectral energy distributions using arcminute-resolution far-ultraviolet (FUV), Hα, far-infrared (FIR), and radio images. Widely-used indicators of the amount of light lost to dust (attenuation) at Hα and in the FUV correlate with each other, although often with substantial scatter. There are two interesting systematic discrepancies. First, Hα attenuations estimated from the Balmer decrement are lower than those estimated from the Hα-to-thermal radio luminosity ratio. Using our full range of data, we tentatively attribute this discrepancy to nonthermal contamination of LMC H ii region radio fluxes. Second, the attenuation at 1500Å and UV spectral slope, β, correlate, although the slope and scatter are substantially different from the correlation first derived for starbursting galaxies by Calzetti et al.

Combining our result with those of Meurer et al. for ultra-luminous infrared galaxies and Calzetti et al. for starbursting galaxies, we conclude that no single relation between β and 1500Å attenuation is applicable to all star-forming systems.

Submitted to: Astrophysical Journal

For preprints, contact ebell@as.arizona.edu
Also available from the URL http://gecko.as.arizona.edu/∼ebell/res.html

Photoionization Models of NGC 346

Mónica Relaño^{1,2}, Manuel Peimbert^{3} and John Beckman^{1,4}

1 Instituto de Astrofísica de Canarias, C/ Vía Láctea s/n, La Laguna, Tenerife, Spain 2 Visiting Astronomer, Instituto de Astronomía, Universidad Nacional Autónoma de México, Apartado Postal 70-264, México 04510, D.F., México. 3 Instituto de Astronomía, Universidad Nacional Autónoma de México, Apartado Postal 70-264, México 04510, D.F., México 4 Consejo Superior de Investigaciones Científicas, Madrid, Spain

We present spherically symmetric and plane parallel photoionization models of NGC 346, an H ii region in the Small Magellanic Cloud. The models are based on Cloudy and on the observations of Peimbert, Peimbert, & Ruiz (2000). We find that approximately 45% of the H ionization photons escape from the H ii region providing an important ionizing source for the low density interstellar medium of the SMC. The predicted I(4363)/I(5007) value is smaller than that observed, probably implying that there is an additional source of energy not taken into account by the models. From the ionization structure of the best model and the observed line intensities we determine the abundances of N, Ne, S, Ar, and Fe relative to O.

Accepted by: The Astrophysical Journal For preprints, contact mpastor@ll.iac.es

Far Ultraviolet Spectroscopic Explorer Observations of a Supernova Remnant in the Line of Sight to HD 5980 in the Small Magellanic Cloud

Charles. G. Hoopes^{1}, Kenneth. R. Sembach^{1}, J. Christopher. Howk^{1}, and William. P. Blair^{1}

1 Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA

We report a detection of far ultraviolet absorption from the supernova remnant SNR 0057 – 7226 in the Small Magellanic Cloud (SMC). The absorption is seen in the Far Ultraviolet Spectroscopic Explorer (FUSE) spectrum of the LBV/WR star HD 5980. Absorption from O vi λ1032 and C iii λ977 is seen at a velocity of +300 km s^{−1} with respect to the Galactic absorption lines, +170 km s^{−1} with respect to the SMC absorption. The O vi λ1038 line is contaminated by H_{2} absorption, but is present. These lines are not seen in the F USE spectrum of Sk80, only ∼ 1′ (∼17 pc) away from HD 5980. No blueshifted O vi λ1032 absorption from the SNR is seen in the F USE spectrum. The O vi λ1032 line in the SNR is well described by a Gaussian with FWHM = 75 km s^{−1}. We find log N(O vi) = 14.33–14.43, which is roughly 50% of the rest of the O vi column in the SMC (excluding the SNR) and greater than the O vi column in the Milky Way halo along this sight line. The N(C iv)/N(O vi) ratio for the SNR absorption is in the range of 0.12–0.17, similar to the value seen in the Milky Way disk, and lower than the halo value, supporting models in which SNRs produce the highly ionized gas close to the plane of the Galaxy, while other mechanisms occur in the halo. The N(C iv)/N(O vi) ratio is also lower than the SMC ratio along this sight line, suggesting that other mechanisms contribute to the creation of the global hot ionized medium in the SMC. The O vi, C iv, and Si iv apparent column density profiles suggest the presence of a multi-phase shell followed by a region of higher temperature gas.

Accepted by: Astrophysical Journal Letters For preprints, contact choopes@pha.jhu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0108049

The Star Formation History of the Large Magellanic Cloud^{1}

Tammy A. Smecker-Hane^{2}, Andrew A. Cole^{3}, John S. Gallagher, III^{4}, and Peter B. Stetson^{5}

1 Based on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract 5-26555. 2 Department of Physics and Astronomy, University of California, Irvine, 4129 Frederick Reines Hall, Irvine, CA 92697– 4575; USA 3 Department of Astronomy, 532A Lederle Grad Research Tower, University of Massachusetts, Amherst, MA 01003, USA 4 Department of Astronomy, University of Wisconsin–Madison, 5534 Sterling Hall, 475 North Charter Street, Madison, WI 53706–1582, USA 5 Dominion Astrophysical Observatory, Herzberg Institute of Astrophysics, 5071 West Saanich Road, Victoria, BC V8X 4M6, Canada

Using WFPC2 aboard the Hubble Space Telescope, we have created deep color-magnitude diagrams in the V and I passbands for approximately 10^{5} stars in a field at the center of the LMC bar and another in the disk. The main–sequence luminosity functions (LFs) from 19 ≤ V ≤ 23.5, the red clump and horizontal branch morphologies, and the differential Hess diagram of the two fields all strongly imply that the disk and bar have significantly different star-formation histories (SFHs). The disk’s SFH has been relatively smooth and continuous over the last ∼ 15 Gyr while the bar’s SFH was dominated by star formation episodes at intermediate ages. Comparison of the LF against predictions based on Padova theoretical stellar evolution models and an assumed age-metallicity relationship allows us to identify the dominant stellar populations in the bar with episodes of star formation that occurred from 4 to 6 and 1 to 2 Gyr ago. These events accounted, respectively, for ∼ 25% and ∼ 15% of its stellar mass. The disk field may share a mild enhancement in SF for the younger episode, and thus we identify the 4 to 6 Gyr episode with the formation of the LMC bar.

Accepted by: The Astrophysical Journal For preprints, contact tsmecker@uci.edu

Also available from the URL http://www.ps.uci.edu/physics/smeckerhane.html

The Distance of the Large Magellanic Cloud Cluster NGC 1866

A.R. Walker^{1}, G. Raimondo^{2,3}, E. Di Carlo^{2}, E. Brocato^{2}, V. Castellani^{4}, V. Hill^{5}

1 Cerro Tololo Inter-American Observatory, NOAO, Casilla 603, La Serena, Chile 2 Osservatorio Astronomico di Collurania, Via M. Maggini, Teramo, Italy 3 Astronomia-Dipartimento di Fisica, Università La Sapienza, P. ls Aldo Moro 2, Rome, Italy 4 Dipartimento di Fisica, Università di Pisa, Pza Torricelli 2, Pisa, Italy 5 European Southern Observatory, Garching, Germany

Hubble Space Telescope V, I photometry of stars in the Large Magellanic Cloud cluster NGC 1866 shows a well defined cluster main sequence down to V=25 mag, with little contamination from field or foreground stars. We use the main sequence fitting procedure to link the distance of NGC 1866 to the Hipparcos determination of the distance for the Hyades MS stars, making use of evolutionary prescriptions to allow for differences in the chemical composition. On this basis we find a true distance modulus for NGC 1866 of 18.35 ± 0.05 mag. If the cluster is assumed to lie in the LMC plane then the LMC modulus is 0.02 mag less.

Accepted by: Astrophysical Journal Letters

Preprints available from the URL http://xxx.lanl.gov/abs/astro-ph/0109147

The optical long ‘Period’ of LMC X-3

C. Brocksopp^{1}, P.J. Groot^{2}, J. Wilms^{3}

1 Astrophysical Research Institute, Liverpool John Moores University, 12 Quays House, Egerton Wharf, Birkenhead CH41 1LD, UK 2 Harvard Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA 3 Institut für Astronomie und Astrophysik, Abt. Astronomie, Waldhäuser Straße 64, D-72076 Tübingen, Germany

We have studied the long term variability of LMC X-3 in optical lightcurves spanning six years, in order to search for optical signatures which could confirm/refute the suggestion that the ‘modulation’ is the result of accretion rate variability rather than accretion disc precession. We find that there is no stable period in the optical lightcurves, that the optical and X-ray lightcurves are correlated, with an X-ray lag and that there is no modulation in the optical colours. We argue that these effects agree with the variable mass accretion rate model proposed by Wilms et al. (2001).

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact cb@astro.livjm.ac.uk

SN 1987A and the Status of Oscillation Solutions to the Solar Neutrino Problem

M. Kachelrieß^{1}, A. Strumia^{1}, R. Tomàs^{2} and J.W.F. Valle^{2}

1 TH Division, CERN, CH–1211 Geneva 23, Switzerland 2 Instituto de Fisica Corpuscular – C.S.I.C. – Universitat de València, E–46071 València, Spain

We study neutrino oscillations and the level-crossing probability P_{LSZ} in power-law potential profiles A(r) ∝ r^{n}. We give local and global adiabaticity conditions valid for all mixing angles ϑ and discuss different representations for P_{LSZ}. For the 1/r^{3} profile typical of supernova envelopes we compare our analytical to numerical results and to earlier approximations used in the literature. We then perform a combined likelihood analysis of the observed SN 1987A neutrino signal and of the latest solar neutrino data, including the recent SNO CC measurement. We find that, unless all relevant supernova parameters (released binding energy, ν̄_{e} and ν̄_{μ,τ} temperatures) are near their lowest values found in simulations, the status of large mixing type solutions deteriorates considerably compared to fits using only solar data. This is sufficient to rule out the vacuum-type solutions for most reasonable choices of astrophysics parameters. The LOW solution may still be acceptable, but becomes worse than the SMA–MSW solution which may, in some cases, be the best combined solution. On the other hand the LMA–MSW solution can easily survive as the best overall solution, although its size is generally reduced when compared to fits to the solar data only.

Submitted to: Physical Review D

For preprints, contact Michael.Kachelriess@cern.ch
Also available from the URL http://your-mirror.arXiv.org/abs/hep-ph/0108100

Thesis Abstract

AGB stars and other red giants in the Magellanic Clouds

M-R.L. Cioni

1 Sterrewacht Leiden, Niels Bohrweg 2, 2333 CA Leiden, The Netherlands

This thesis focuses on the analysis and interpretation of observations of the whole Magellanic Clouds obtained as part of the Deep Near–Infrared Southern Sky Survey (DENIS – Epchtein et al. 1997). Note that very few of the earlier measurements were surveys (panoramic view of the whole system). The compilation of a catalogue with I, J and K_{S} data for 1.3 and 0.3 million sources towards the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC), respectively, has allowed us to investigate several important issues: the distance, the morphology, the parameters of the LMC disk structure, the spectral type and the variability of AGB stars.

PhD Thesis to be defended at the Leiden University on September 20th under the direction of Harm J. Habing. For copies or inquiries, contact mrcioni@strw.leidenuniv.nl

Abstracts of Non-Refereed Papers

High-velocity clouds: a diverse phenomenon

B.P. Wakker^{1}

1 University of Wisconsin, 475 N. Charter St., Madison, WI53706, USA

In this contribution the current state of knowledge about the high-velocity clouds (HVCs) is summarized. Recent progress has shown that the HVCs are a diverse phenomenon. The intermediatevelocity clouds (IVCs) are likely to be part of a Galactic Fountain. The Magellanic Stream is a tidal remnant. HVC complex C (possibly complexes A and GCN) are low-metallicity clouds near the Galaxy; they could be remnants of the formation of the Galaxy or old tidal streams extracted from nearby dwarf galaxies. Having a substantial number of H I HVCs dispersed throughout the Local Group seems incompatible with the observed H I mass function of galaxies. Finally, FUSE finds highvelocity O VI, some of which is clearly associated with H I HVCs, but some which is not.

To appear in: 17th IAP Colloquium: Gaseous Matter in Galaxies and Intergalactic Space

For preprints, contact wakker@astro.wisc.edu

The small-scale structure of the Magellanic Stream

S. Stanimirovic^{1}, J. M. Dickey^{2} and A. M. Brooks^{3}

1 Arecibo Observatory, NAIC/Cornell University, HC 3 Box 53995, Arecibo, Puerto Rico 00612, USA 2 Astronomy Department, University of Minnesota, 116 Church Street SE, Minneapolis, MN 55455, USA 3 Macalester College, Saint Paul, MN 55105, USA

We have mapped in neutral hydrogen (HI) two regions at the northern tip of the Magellanic Stream, known as MS V and MS VI, using the Arecibo telescope. The small-scale structure of the MS shows clumpy and head-tail morphology. The spatial power spectrum of this star-free intergalactic medium has a power-law behavior with the density slope of −3.8. A gradual steepening of the power-law slope is seen when increasing the thickness of velocity slices.

To appear in: Proceedings of the Meeting on Gaseous Matter in Galactic and Intergalactic Space, XVIIth IAP Colloquium For preprints, contact sstanimi@naic.edu

Also available from the URL http://arXiv.org/abs/astro-ph/0109211

The youngest massive star clusters in the Magellanic Clouds

M. Heydari-Malayeri^{1}, V. Charmandaris^{2}, L. Deharveng^{3}, M.R. Rosa^{4}, D. Schaerer^{5}, & H. Zinnecker^{6}

1Observatoire de Paris, DEMIRM, 61 Ave de l’Observatoire, F–75014 Paris, France 2Astronomy Department, Cornell Univ., Ithaca, NY 14853, USA 3Observatoire de Marseille, 2 Place Le Verrier, F-13248 Marseille Cedex 4, France 4STECF - European Southern Observatory, D-85748 Garching, Germany 5Observatoire Midi-Pyrenees, 14, Avenue E. Belin, F-31400 Toulouse, France 6Astrophysikalisches Institut Potsdam, D-14482 Potsdam, Germany

High resolution observations with HST have recently allowed us to resolve and study several very tight clusters of newly born massive stars in the Magellanic Clouds. Situated in an extremely rare category of H ii regions, being only 5′′ to 10′′ across and of high excitation and extinction, these stars are just hatching from their natal molecular clouds. Since the SMC is the most metal-poor galaxy observable with very high angular resolution, this work may provide valuable templates for addressing issues of star formation in the very distant metal-poor galaxies of the early Universe.

To appear in: Conference proceedings of the XVIIth IAP Colloquium, “Gaseous matter in Galaxies and Intergalactic Space”, Paris France, June 2001

For preprints, contact mohammad.heydari-malayeri@obspm.fr
Also available from the URL http://arXiv.org/abs/astro-ph/0109034

Mass Segregation in the Young SMC Clusters NGC 330

M.Sirianni^{1}, A. Nota^{2,3}, G. De Marchi^{2,3}, C. Leitherer^{2} and M. Clampin^{2}

1 Johns Hopkins University, 3400 N Charles Street, Baltimore, MD 21218, USA 2 STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA 3 Affiliated with the Astrophysics Division, Space Science Department of the European Space Agency.

We present a new study of the low end of the stellar IMF of NGC 330, the richest young star cluster in the Small Magellanic Cloud (SMC). Using deep broadband images taken with the HST/WFPC2 we have derived the cluster’s luminosity function and constructed the Initial Mass Function (IMF) in the mass range 1 − 7 M_{⊙}. We have investigated the IMF as a function of the radial distance from the cluster center. We find that, after correction for background contamination, the IMF is fairly homogeneous with a slope slightly steeper than Salpeter’s in the central regions of the cluster (< 40′′) but becomes increasingly steeper with distance, indicating a preponderance of massive stars in the core of the cluster. NGC 330 is one of the first clusters for which evidence of mass segregation is directly found.

To appear in: ”Extragalactic Stars Clusters” IAU Symp. 207, 2001 Eva K. Grebel, Doug Geisler and Dante Minniti, eds.

For preprints, contact sirianni@pha.jhu.edu